OpenStudy (anonymous):

what is 78631-452672

OpenStudy (across):

Is this a joke? I'm sorry, but I'd rather be honest.

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=78631-452672

OpenStudy (anonymous):

huh, why would it be a joke

OpenStudy (across):

I believe we're here to provide insight and help others understand ther problems. What insight is there to provide here?

OpenStudy (anonymous):

i am not kiding what is it

OpenStudy (anonymous):

well things like borrowing 10 from left

OpenStudy (anonymous):

where are you having difficulty with it ?

OpenStudy (anonymous):

help!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!!

OpenStudy (anonymous):

please

OpenStudy (anonymous):

what kind of help

OpenStudy (across):

The asker is using a computer to ask questions in an online-based, live-feedback system. If he or she can do that, then why can't he or she not know how to subtract two numbers?

OpenStudy (anonymous):

^^^Do that

OpenStudy (anonymous):

78631-452672 subtract 73631 from 452672 and put a negative sign before the answer

OpenStudy (amistre64):

i think its modular :)

myininaya (myininaya):

maybe they are trying to find the difference without a calculator some people actually don't know how to do that yet

OpenStudy (amistre64):

maybe, but highly unlikely ....

OpenStudy (anonymous):

^ yeah, that's better

OpenStudy (anonymous):

7 viewing .....where is the Questioner ?

OpenStudy (anonymous):

looking for a calculator

myininaya (myininaya):

(14) (3)(15)(12) 4 5 2 6 7 2 - 7 8 6 3 1 ------------- 3 7 4 0 4 1 so 78631-452672=-374041

OpenStudy (anonymous):

I wouldn't like it when I post a question and people made comment like, "too easy" ,

OpenStudy (amistre64):

78631- 452672 = N 78631 = N + 452672 -70000 -70000 --------------------- 8631 = N + 382672 -8000 -8000 --------------------- 631 = N + 374672 -600 -600 --------------------- 31 = N + 374072 -30 -30 -------------------- 1 = N + 374042 -1 -1 ------------------ 0 = N + 374041 -374041 -374041 ---------------------- -374041 = N